# Self-ReS: Self-Reflection in Large Vision-Language Models for Long Video Understanding

### v1

@@ Abstract
Large Vision-Language Models (LVLMs) demonstrate remarkable performance in short-video tasks such as video question answering, but struggle in long-video understanding. The linear frame sampling strategy, conventionally used by LVLMs, fails to account for the non-linear distribution of key events in video data, often introducing redundant or irrelevant information in longer contexts while risking the omission of critical events in shorter ones. To address this, we propose Self-ReS, a non-linear spatiotemporal self-reflective sampling method that dynamically selects key video fragments based on user prompts. Unlike prior approaches, Self-ReS leverages the inherently sparse attention maps of LVLMs to define reflection tokens, enabling relevance-aware token selection without requiring additional training or external modules. Experiments demonstrate that Self-ReS can be seamlessly integrated into strong base LVLMs, improving long-video task accuracy and achieving up to 46% faster inference speed within the same GPU memory budget. The code can be found here {hidden for submission}.
@@ III-A Spatiotemporal Self-Reflection
$\text{SRA}(q_{l},K_{V_{l}})=\frac{q_{l}\cdot K_{V_{l}}}{\sqrt{d_{k}}}.$ (2)
@@ III-B Converging Reflection Paths
Regardless of the chosen strategy, Self-ReS converges into a unified new inference path sequence. This sequence is forwarded using vanilla attention and reassigned positional embeddings. This path focuses exclusively on the most salient visual tokens, identified through the LLM’s self-reflective capabilities guided by the user query. By leveraging these tokens, the final model prediction effectively captures global inter-segment dependencies, enabling reasoning over a larger context of the input video. Notably, this is achieved without actually increasing the model’s default context size or memory requirements, ensuring efficient handling of non-linear video events.
@@ III-C Complexity of Self-ReS
Our proposed SRA effectively computes saliency scores for visual tokens, similar to regular attention mechanisms but with reduced computational overhead. In standard LVLM inference, the computational cost of attention is quadratic over the total number of input tokens - $O(N^{2})$, where $N=N_{r}*R$. In Self-ReS, while attention computation remains quadratic, it is computed over a much smaller number of input tokens for each reflection path: $O(R\times N_{r}^{2})=O(N_{r}^{2})$. Thus, attention computation is reduced by a factor of $R^{2}$. SRA (Eq. 2) has linear complexity $O(N_{v})$, which is then applied to each reflection path $O(R\cdot N_{V})$. Therefore, it does not affect the overall complexity of the LVLM inference flow.
@@ IV-C Long Video Understanding
Large VideoLLaMA2[23] Q-Former 32 57.6 62.4 45.6
@@
Mid-sized MovieChat[9] Memory 2048 - - 18.0
@@
LLaVA-Video Pooling 32 52.2 62.5 66.5
@@
LLaVA-Video[3] Self-ReS 128 (32) 54.6 64.3 67.9

### v2

# João Pereira1,2, Vasco Lopes2,3, David Semedo1, João Neves3

@@ Abstract
Large Vision-Language Models (LVLMs) demonstrate remarkable performance in short-video tasks such as video question answering, but struggle in long-video understanding. The linear frame sampling strategy, conventionally used by LVLMs, fails to account for the non-linear distribution of key events in video data, often introducing redundant or irrelevant information in longer contexts while risking the omission of critical events in shorter ones. To address this, we propose Self-ReS, a non-linear spatiotemporal self-reflective sampling method that dynamically selects key video fragments based on user prompts. Unlike prior approaches, Self-ReS leverages the inherently sparse attention maps of LVLMs to define reflection tokens, enabling relevance-aware token selection without requiring additional training or external modules. Experiments demonstrate that Self-ReS can be seamlessly integrated into strong base LVLMs, improving long-video task accuracy and achieving up to 46% faster inference speed within the same GPU memory budget. The code can be found here: https://github.com/jaca-pereira/Self-ReS.
@@ III-A Spatiotemporal Self-Reflection
$\text{SRA}(q_{l},K_{V_{l}})=\frac{q_{l}\cdot K^{T}_{V_{l}}}{\sqrt{d_{k}}}.$ (2)
@@ III-B Converging Reflection Paths
Regardless of the chosen strategy, Self-ReS converges into a unified new inference path sequence. This sequence is forwarded using vanilla attention and reassigned positional embeddings (see section -A of the Appendix). This path focuses exclusively on the most salient visual tokens, identified through the LLM’s self-reflective capabilities guided by the user query. By leveraging these tokens, the final model prediction effectively captures global inter-segment dependencies, enabling reasoning over a larger context of the input video. Notably, this is achieved without actually increasing the model’s default context size or memory requirements, ensuring efficient handling of non-linear video events.
@@ III-C Complexity of Self-ReS
Our proposed SRA effectively computes saliency scores for visual tokens, similar to regular attention mechanisms but with reduced computational overhead. In our baseline LVLM, the computational cost of attention is quadratic over the total number of input tokens - $O(R*N_{r}^{2})$, where $R$ is the number of reflection paths and $N_{r}$ is the number of input tokens per reflection path. In Self-ReS, while attention computation remains quadratic, it is computed over a smaller number of input tokens split between reflection paths: $O(R*N_{r}^{2})=O(R*(N/R)^{2})=O(N^{2}/R)$. Thus, while Self-ReS maintains quadratic complexity, it is reduced by a factor of $R$. SRA (Eq. 2) has linear complexity $O(N_{v})$, which is then applied to each reflection path $O(R\cdot N_{V})$. Therefore, it does not affect the overall complexity of the LVLM inference flow.
@@ IV-C Long Video Understanding
Large >=>=> =72B VideoLLaMA2[23] Q-Former 32 57.6 62.4 45.6
@@
Mid-Sized 7B MovieChat[9] Memory 2048 - - 18.0
@@
LLaVA-Video[3] Pooling 32 52.2 62.5 66.5
@@
LLaVA-Video Self-ReS 128 (32) 54.6 64.3 67.9